OpenStudy (anonymous):

Find the equation of the normal to y=log base e (x+2) which is parallel to the line with equation y+3x-5=0

OpenStudy (anonymous):

@kropot72

OpenStudy (anonymous):

this is how i did it

OpenStudy (anonymous):

y+3x-5=0 y=-3x+5 gradient = -3 find the derivative of y=log base e (x+2): y'=1/(x+2) therefore derivative of y is equal to the gradient: 1/(x+2)=-3 x=-7/3

OpenStudy (anonymous):

am i solving it correctly? @kropot72

OpenStudy (anonymous):

the problem is i get an error for solving for y when i plug in the value of x=-7/3

OpenStudy (anonymous):

do you think there is something wrong in the question?

OpenStudy (anonymous):

help @ganeshie8

OpenStudy (dumbcow):

you set the tangent line equal to gradient, not the normal

OpenStudy (anonymous):

ok

OpenStudy (dumbcow):

the slope of tangent is always positive on a log function anyway, the normal slope is the opposite reciprocal to the tangent --> -(x+2) = -3

OpenStudy (anonymous):

so -1/m is the gradient of the normal

OpenStudy (dumbcow):

correct

OpenStudy (anonymous):

ok now i get it

OpenStudy (anonymous):

so m in this case is -1/3

OpenStudy (kropot72):

The product of the slope of y = log base e (x + 2) and -3 must equal -1. \[\Large \frac{1}{x+2} \times -3 = -1\] Solve for x. Then plug the value of x into y = log base e (x +2). The resulting y value is then used to find the required equation , which is of the form y = -3x + b

OpenStudy (anonymous):

will it become -1/3=1/(x+2)???

OpenStudy (dumbcow):

@Chad123 no m = -3 (slope of parallel line) gradient of normal = -1/y'

OpenStudy (kropot72):

\[\Large \frac{1}{x+2}=\frac{1}{3}\]

OpenStudy (anonymous):

gradient of normal then is = -1/(1/(x+2))

OpenStudy (anonymous):

am i correct?

OpenStudy (kropot72):

Surely the gradient of the normal must equal -3.

OpenStudy (anonymous):

can someone please show me how it is done step by step because im kind of lost

OpenStudy (dumbcow):

follow @kropot72 posts

OpenStudy (kropot72):

The value of x when the normal is parallel to y + 3x - 5 = 0 is 1. Put x = 1 into y = log base e (x + 2) to get y = ln 3 = 1.099. Plug y = 1.099 and x = 1 into y = -3x + b and find the value of b. Then you will have the solution.

OpenStudy (anonymous):

where did x=1 come from?

OpenStudy (kropot72):

The solution to: \[\Large \frac{1}{x+2}=\frac{1}{3}\]

OpenStudy (anonymous):

1/3 came from plugging m=-3 into -1/m yielding 1/3

OpenStudy (anonymous):

am i right?

OpenStudy (kropot72):

Looks right to me, although I didn't put it that way.

OpenStudy (anonymous):

solving for x gives 1

OpenStudy (anonymous):

now it makes sense

OpenStudy (anonymous):

thank you so much :)

OpenStudy (kropot72):

Gr8! You're welcome :)

OpenStudy (anonymous):

ill still have to solve the equation of the normal which i would probably know how to solve

OpenStudy (kropot72):

Just do exactly as my previous post says: "The value of x when the normal is parallel to y + 3x - 5 = 0 is 1. Put x = 1 into y = log base e (x + 2) to get y = ln 3 = 1.099. Plug y = 1.099 and x = 1 into y = -3x + b and find the value of b. Then you will have the solution." Note: We already know that the form of the equation of the normal is: y = -3x + b

OpenStudy (kropot72):

So you just need to find the value of b by substitution of x = 1 and y = 1.099.

OpenStudy (anonymous):

im using the equation of the normal formula y-f(a)=(-1/f'(a)) (x-a)

OpenStudy (kropot72):

You don't need to use that equation at this point. Please just do the substitution as I described.

OpenStudy (kropot72):

ust plug x = 1 and y = 1.099 into y = -3x + b. Then you will have the value of b, and the equation of the normal

OpenStudy (anonymous):

i will

OpenStudy (anonymous):

thank u :)

OpenStudy (anonymous):

you were very helpful

OpenStudy (kropot72):

np